OpenStudy (anonymous):

The sides of a rhombus abcd are parallel to the lines y=x+2'and y=7x+3 if the diagonals of the rhombus intersect at point o(1,2)

OpenStudy (anonymous):

y=x+2 and y=7x+3 x+2=7x+3 x=-1/6 y=7(-1/6)+3 y=11/6